OpenStudy (anonymous):

Solve the equation: \[(2+\frac{ 3 }{ 2 })^2-(2+\frac{3 }{ x })-12=0\]

OpenStudy (anonymous):

@dumbcow you mind helping me with this?

OpenStudy (anonymous):

@dumbcow now any hints on how to solve this one? lol

OpenStudy (dumbcow):

try to isolate the "3/x" 2+ 3/2 = 7/2

OpenStudy (dumbcow):

\[(\frac{7}{2})^2 -12-2 = \frac{3}{x}\]

OpenStudy (anonymous):

so then its \[\frac{ 49 }{ 4 }-12-2=\frac{ 3 }{ x }\]

OpenStudy (anonymous):

which equals \[\frac{ -7 }{ 4 }=\frac{ 3 }{ x }\]

OpenStudy (dumbcow):

correct ... cross-multiply and solve for x

OpenStudy (anonymous):

x=-12/7?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

alright thanks man

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

and i just noticed im an idiot -_-

OpenStudy (anonymous):

lol the beginning (2+3/2)^2 is (2+3/x)^2

OpenStudy (anonymous):

\[(2+\frac{ 3 }{ x })^2-(2+\frac{ 3 }{ x })-12=0\]

OpenStudy (dumbcow):

yeah i was afraid of that... i should have said something ok its not hard just make a substitution let u = 2 + 3/x --> u^2 - u - 12 = 0 solve for u